OpenStudy (anonymous):

Simplify the expression (2x − 6)(3x2 − 3x − 6).

OpenStudy (cwtan):

factorise 3x2-3x-6 it will be 3(x2-x-2) =3(x-2)(x+1) 2x-6=2(x-3) so combine 6(x-3)(x-2)(x+1)

OpenStudy (anonymous):

so its 6x^3 − 24x^2 + 6x − 36?

OpenStudy (cwtan):

6x^3 − 24x^2 + 6x "+"36

OpenStudy (anonymous):

expand the equation...u will get the answer;) (2x-6)(3x^2-3x-6) = 6x^3-6x^2-12x-18x^2+18x+36 6x^3-24x^2+6x+36 simplify the answer.. n u will get x^3-4x^2+x+6